OpenStudy (anonymous):

What are the steps for multiplying monomial algebraic terms where they have exponents with like bases?

OpenStudy (campbell_st):

add the powers \[3x^a \times 2x^b = (3 \times 2)x^{a + b} = 6x^{a + b}\]